A sample of C3H8 has 1.60×1024 H atoms.

How many carbon atoms does the sample contain?

number of carbon atoms:

?×10 ? C atoms

What is the total mass of the sample?

mass: ? ×10 ? g C3H8
Answer #1

1)

1 C3H8 has 3 Carbon atoms and 8 H atoms.

So,

number of C = (3/8)*number of H atoms

= (3/8)*(1.60*10^24)

= 6.00*10^23 atoms

Answer: 6.00*10^23 C atoms

2)

number of C3H8 molecule = 1/8 * number of H atoms

= (1/8)*1.60*10^24

= 2.0*10^23

use:

number of mol = number of molecules / Avogadro’s number

number of mol = 2.0*10^23 / 6.022*10^23 mol

number of mol = 0.3321 mol

Molar mass of C3H8,

MM = 3*MM(C) + 8*MM(H)

= 3*12.01 + 8*1.008

= 44.094 g/mol

use:

mass of C3H8,

m = number of mol * molar mass

= 0.3321 mol * 44.09 g/mol

= 14.64 g

Answer: 1.46*10^1 g